# Relational Representation Distillation

### v3

@@ 5 Conclusions
Our method offers a significant advancement in KD by maintaining relational consistency between teacher and student models. RRD leverages a large memory buffer of teacher samples to align their output distributions, ensuring consistent relational structures throughout the learning process. Unlike traditional approaches, RRD uses pairwise similarities and a relaxed contrastive loss to enhance the robustness and performance of the student model without explicit negative instances. Through comprehensive testing on CIFAR-100, ImageNet, TIN-200, and STL-10 datasets, RRD consistently outperforms traditional KD techniques and state-of-the-art methods, demonstrating substantial improvements in accuracy, robustness, and transferability of learned representations.

### v4

@@ 5 Conclusions
Our method advances KD by maintaining relational consistency between teacher and student models. RRD leverages a large memory buffer of teacher samples to align output distributions, ensuring consistent relational structures throughout learning. Unlike traditional approaches, RRD uses pairwise similarities and a relaxed contrastive loss to improve robustness and performance of the student model without explicit negatives. Extensive experiments on CIFAR-100, ImageNet, STL-10, and Tiny ImageNet show that RRD consistently outperforms state-of-the-art KD methods, demonstrating improved accuracy and transferability of representations.